# Stabilizer Rényi Entropy and its Transition in the Coupled Sachdev-Ye-Kitaev Model

### v2

@@ Abstract
In this Letter, we establish a general path-integral formalism for calculating the SRE in a class of solvable Sachdev-Ye-Kitaev (SYK) models [66, 67, 68], which allows for exact analysis in the large-$N$ limit. We focus on the thermal ensemble of the Maldacena-Qi coupled SYK model [69, 70, 71] (illustrated in FIG. 1), which exhibits a Hawking-Page transition from a low-temperature wormhole phase to a high-temperature black hole phase. We show that its SRE, while remaining extensive as the inverse temperature $\beta$ is tuned, exhibits three first-order transitions. Notably, one of these transitions cannot be detected by the system’s equilibrium properties and represents an intrinsic transition of the SRE [49]. We explain that this transition arises from a change in connectivity between different branches of the path-integral contour, closely analogous to the celebrated replica wormholes [72, 73, 74]. Our results pave the way for studying the SRE of strongly interacting fermions using large-$N$ solvable models and demonstrate that it can reveal hidden structures in many-body systems that are inaccessible to equilibrium observables, with potential connections to holography.
@@
The distribution of $c_{\bm{v}}(\rho)$ encodes information about the quantum magic of the state. For pure stabilizer states, $c_{\bm{v}}(\rho)=\pm 1$ if $\Psi_{\bm{v}}$ is a stabilizer, and $c_{\bm{v}}(\rho)=0$ otherwise. This observation motivates the introduction of the SRE $\tilde{M}_{2}(\rho)=M_{2}(\rho)-S_{2}(\rho)$ in Refs. [34, 36, 35], with
@@
Therefore, evaluating the SRE amounts to inserting operators on each site of the replicated system. When the density matrix is represented as an imaginary-time path integral [77], these operators are inserted at a single imaginary time, chosen as $\tau=0$, as sketched in FIG. 2.
@@
Next, we introduce a weak coupling between the two sides with $\mu/J=0.1$. In the high-temperature limit, the coupling effect is perturbative, and $M_{2}(\beta)\approx N\big(\ln 2-O(\mu^{2}\beta^{2})\big)$. As the inverse temperature increases, we observe a first-order transition of $S_{\text{SRE}}$ near $\beta^{*}_{\text{SRE}}J\approx 19$, evidenced by the coexistence of two distinct solutions to Eq.(10), obtained from different initialization conditions in the iteration. The physical solution is the one with smaller $M_{2}(\beta)$ and the SRE becomes singular. Similar singularities have been observed in quantum Monte Carlo simulations of quantum spin models [49]. Fig.3(b) shows the Green’s function $G^{(11)}_{LL}(\tau,\tau^{\prime})$. The solution corresponding to the high-temperature saddle resembles the thermal Green’s function, approaching $1/2$ at $\tau=0^{+}$ and $\tau^{\prime}=\beta^{-}$. This behavior indicates that the SWAP operators are effectively disrupted by the fluctuating field $\sigma$, leading to a disconnection between replicas. In contrast, the Green’s function at low temperatures decays as $\tau-\tau^{\prime}$ increases, consistent with full connectivity between replicas (1) and (2), where $\tau=0^{+}$ and $\tau^{\prime}=\beta^{-}$ correspond to maximal imaginary-time separation. Therefore, the transition reflects a qualitative change in replica connectivity. If we further increase $\beta$, another first-order transition appears near $\beta^{*}_{\text{HP}}J\approx 27$. This is the Hawking-Page transition of the thermal partition function, originating from the contribution of $S_{\beta}$ in the relation $M_{2}(\rho)=S_{\text{SRE}}-4S_{\beta}+N\ln 2$. The physical solution corresponds to the branch with smaller $S_{\beta}$, and therefore larger $M_{2}(\beta)$. For $\beta J\rightarrow\infty$, $M_{2}(\beta)$ saturates to a value significantly smaller than in the case with $\mu=0$. In the supplementary material [78], we further show that $M_{2}(\infty)=N(2\ln 2-4s_{0})<N\ln 2$ for $\mu/J\rightarrow 0$, where $s_{0}\approx 0.2324$ is the zero-temperature entropy density of the SYK4 model [68]. Consequently, $M_{2}(\infty)$ is discontinuous at $\mu=0$ as $\mu/J$ is varied. For larger $\mu/J=0.2$, the transitions of $M_{2}(\beta)$ turn into crossovers, while the qualitative behavior remains unchanged.
@@
Acknowledgement. We thank Langxuan Chen, Yingfei Gu, Haolin Jiang, Zeyu Liu, Yuzhi Tong, and Zhao-Yi Zeng for helpful discussions. This project is supported by the NSFC under grant 12374477, the Shanghai Rising-Star Program under grant number 24QA2700300, and the Innovation Program for Quantum Science and Technology 2024ZD0300101.

### v3

@@ Abstract
In this Letter, we establish a general path-integral formalism for calculating the SRE in a class of solvable Sachdev-Ye-Kitaev (SYK) models [66, 67, 68, 69], which allows for exact analysis in the large-$N$ limit. We focus on the thermal ensemble of the Maldacena-Qi coupled SYK model [70, 71, 72] (illustrated in FIG. 1), which exhibits a Hawking-Page transition from a low-temperature wormhole phase to a high-temperature black hole phase. We show that its SRE, while remaining extensive as the inverse temperature $\beta$ is tuned, exhibits three first-order transitions. Notably, one of these transitions cannot be detected by the system’s equilibrium properties and represents an intrinsic transition of the SRE [49]. We explain that this transition arises from a change in connectivity between different branches of the path-integral contour, closely analogous to the celebrated replica wormholes [73, 74, 75]. Our results pave the way for studying the SRE of strongly interacting fermions using large-$N$ solvable models and demonstrate that it can reveal hidden structures in many-body systems that are inaccessible to equilibrium observables, with potential connections to holography.
@@
A special class of quantum states, known as stabilizer states, admits an efficient classical representation. In fermionic systems [40], each pure stabilizer state is uniquely specified as the simultaneous +1 eigenstate of $N$ independent, mutually commuting Majorana strings $\{\pm\Psi_{\bm{v}_{1}},\pm\Psi_{\bm{v}_{2}},...,\pm\Psi_{\bm{v}_{N}}\}$, which serve as the generators of the stabilizer group. Consequently, $c_{\bm{v}}(\rho)=\pm 1$ if $\Psi_{\bm{v}}$ is an element of the stabilizer group, and $c_{\bm{v}}(\rho)=0$ otherwise. This observation motivates the introduction of the SRE $\tilde{M}_{2}(\rho)=M_{2}(\rho)-S_{2}(\rho)$ in Refs. [34, 36, 35], to quantify deviations from stabilizer states, i.e., the presence of quantum magic, with
@@
Therefore, evaluating the SRE amounts to inserting operators on each site of the replicated system. When the density matrix is represented as an imaginary-time path integral [78], these operators are inserted at a single imaginary time, chosen as $\tau=0$. A sketch is presented in Fig. 2, where each copy of the imaginary-time evolution $e^{-\beta H}$ is depicted as a closed circle, with different copies shown in distinct colors. The black dots denote operator insertions that couple the different copies. A similar replica trick has been widely used in calculations of entanglement entropy [79].
@@
Next, we introduce a weak coupling between the two sides with $\mu/J=0.1$. In the high-temperature limit, the coupling effect is perturbative, and $M_{2}(\beta)\approx N\big(\ln 2-O(\mu^{2}\beta^{2})\big)$. As the inverse temperature increases, we observe a first-order transition of $S_{\text{SRE}}$ near $\beta^{*}_{\text{SRE}}J\approx 19$, evidenced by the coexistence of two distinct solutions to Eq.(10), obtained from different initialization conditions in the iteration. The physical solution is the one with smaller $M_{2}(\beta)$ and the SRE becomes singular. Similar singularities have been observed in quantum Monte Carlo simulations of quantum spin models [49]. Fig.3(b) shows the Green’s function $G^{(11)}_{LL}(\tau,\tau^{\prime})$. The solution corresponding to the high-temperature saddle resembles the thermal Green’s function, approaching $1/2$ at $\tau=0^{+}$ and $\tau^{\prime}=\beta^{-}$. This behavior indicates that the SWAP operators are effectively disrupted by the fluctuating field $\sigma$, leading to a disconnection between replicas. In contrast, the Green’s function at low temperatures decays as $\tau-\tau^{\prime}$ increases, consistent with full connectivity between replicas (1) and (2), where $\tau=0^{+}$ and $\tau^{\prime}=\beta^{-}$ correspond to maximal imaginary-time separation. Therefore, the transition reflects a qualitative change in replica connectivity. From the perspective of fermionic modes, the coupling between the two sides becomes dominant in the low-temperature saddle. This drives the system toward a target stabilizer state with stabilizers $\{-i\psi_{j,L}\psi_{j,R}\}$, thereby leading to a reduction in $M_{2}(\beta)$.
If we further increase $\beta$, another first-order transition appears near $\beta^{*}_{\text{HP}}J\approx 27$. This is the Hawking-Page transition of the thermal partition function, originating from the contribution of $S_{\beta}$ in the relation $M_{2}(\rho)=S_{\text{SRE}}-4S_{\beta}+N\ln 2$. The physical solution corresponds to the branch with smaller $S_{\beta}$, and therefore larger $M_{2}(\beta)$. For $\beta J\rightarrow\infty$, $M_{2}(\beta)$ saturates to a value significantly smaller than in the case with $\mu=0$. In the supplementary material [80], we further show that $M_{2}(\infty)=N(2\ln 2-4s_{0})<N\ln 2$ for $\mu/J\rightarrow 0$, where $s_{0}\approx 0.2324$ is the zero-temperature entropy density of the SYK4 model [68]. Consequently, $M_{2}(\infty)$ is discontinuous at $\mu=0$ as $\mu/J$ is varied. For larger $\mu/J=0.2$, the transitions of $M_{2}(\beta)$ turn into crossovers, while the qualitative behavior remains unchanged.
@@
Acknowledgement. We thank Langxuan Chen, Yingfei Gu, Haolin Jiang, Zeyu Liu, Yuzhi Tong, and Zhao-Yi Zeng for helpful discussions. This project is supported by the NSFC under grant 12374477, the Shanghai Rising-Star Program under grant number 24QA2700300, and the Quantum Science and Technology-National Science and Technology Major Project 2024ZD0300101.